# Multi-Agent Collaborative Data Selection for Efficient LLM Pretraining

### v1

@@ A.2.1 Details of Pretraining Models Architecture
Hyperparameter 1.3B Model Value 370M Model Value
Vocabulary Size 32,000 32,000
MLP Ratio 8/3 8/3
Hidden Dimension Size 2048 1024
Number of Layers 24 24
Number of Attention Heads 16 8
Number of KV Attention Heads 16 8
RoPE Base 10,000 10,000
Maximum Context Window Length 1024 1024
Number of Parameters 1,345,423,360 (1.3B) 373,867,520 (370M)
A.2.2 Details of Baseline Method Implementation
@@
A.2.3 Details Agent Weight Initialization
We employ an agent weight initialization technique within the RegMix (Liu et al., 2024) framework, which is crucial for the effective training of proxy models. Our dataset is organized into three distinct categories: domain, quality, and topic. For each category, we initialize the data weights based on the original proportions across 512 configurations and subsequently train a TinyLlama-1M with 1 billion tokens as a proxy model for each configuration. We evaluate this model on previously unseen data mixtures, specifically using validation set loss, following RegMix, for assessment. We then fit a regression model based on the performance results of the 512 proxy models to predict the optimal data mixture for training large-scale LLMs. The results of the linear regression analysis and Spearman correlation of the loss prediction performance are presented in 7.
@@ A.2.3 Details Agent Weight Initialization
Figure 7: We present the results of the linear regression analysis and Spearman correlation regarding the loss prediction performance and the weights of each candidate data-(a) Topic, (b) Quality Interval and (c) Domain after mixture across all categories.
@@
A.3 Full Experimental Results
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random sample Uniform-30B 54.3 23.4 22.3 23.9 18.6 39.8 52.8 19.2 55.4 30.0 34.0 Uniform-60B 55.2 24.6 22.5 23.4 21.0 39.7 51.9 19.5 59.8 33.1 35.1 Perplexity-based data selection PPL 49.3 20.1 22.4 23.6 16.2 36.0 48.1 18.8 61.4 29.3 32.5 Classifier-based data selection QuRating-Facts 56.1 23.3 22.4 24.8 21.6 39.2 54.1 19.9 61.5 31.6 35.5 QuRating-Req 54.9 24.4 23.2 25.2 21.4 38.1 54.5 20.6 61.6 31.3 35.5 QuRating-Writing 53.6 23.2 23.4 23.2 21.0 38.1 52.8 19.7 59.4 31.6 34.6 QuRating-Edu 57.0 24.4 22.0 25.0 20.4 40.3 53.7 20.2 60.1 32.2 35.5 FineWeb-Edu 53.8 23.4 21.8 23.9 19.8 39.2 51.7 20.8 59.7 32.0 34.6 DSIR-Book 45.4 20.8 22.0 23.0 18.8 39.9 54.6 19.7 58.3 30.8 33.3 DSIR-Wiki 50.6 21.1 21.6 23.0 19.2 36.6 53.0 19.8 60.5 29.2 33.5 Domain mixing methods DOGE 49.4 21.8 22.5 23.0 18.0 38.0 52.7 19.9 60.0 30.0 33.5 DoReMi 50.1 20.2 22.5 23.7 17.8 38.7 52.8 19.7 58.6 30.8 33.5 DMLaw 49.6 21.9 23.2 23.6 17.8 38.6 51.8 20.1 60.4 29.0 33.6 RegMix 50.0 22.3 22.1 22.9 18.8 38.0 52.8 19.9 58.9 31.2 33.7 Influence functions MATES 50.0 21.4 22.7 25.3 19.0 39.8 53.6 21.3 59.9 32.1 34.5 Multi-Agent Collaboration (Ours) 61.1 28.2 22.6 26.0 24.4 38.2 54.2 19.5 61.0 29.8 36.5
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random sample Uniform-30B 54.6 23.0 22.1 24.9 18.8 40.3 52.9 21.5 53.0 29.8 34.1 Uniform-60B 58.8 25.5 23.0 27.2 20.0 41.8 53.6 19.6 56.9 32.7 35.9 Perplexity-based data selection PPL 50.6 21.3 22.7 25.2 15.6 37.7 48.9 20.1 61.5 22.3 32.6 Classifier-based data selection QuRating-Facts 59.5 25.7 22.6 25.9 19.8 40.2 54.6 19.2 60.8 24.8 35.3 QuRating-Req 59.3 25.9 22.7 26.1 19.6 39.7 53.7 20.5 58.5 22.7 34.9 QuRating-Writing 56.9 25.7 23.1 26.0 20.4 41.1 53.6 20.2 51.4 22.6 34.1 QuRating-Edu 60.8 26.5 22.5 26.7 20.2 41.4 54.6 20.6 55.5 22.7 35.1 FineWeb-Edu 56.2 25.7 22.3 26.2 20.6 40.1 50.5 19.7 56.6 31.4 34.9 DSIR-Book 48.7 21.0 22.6 25.6 18.6 42.5 53.7 19.5 57.9 22.9 33.3 DSIR-Wiki 53.2 22.4 22.6 25.3 17.6 37.1 52.7 21.4 61.6 24.2 33.8 Domain mixing methods DOGE 52.4 21.9 22.4 27.0 17.4 39.9 52.0 18.2 57.8 29.8 33.9 DoReMi 53.2 21.4 22.2 24.7 18.2 38.4 50.9 20.6 59.7 31.1 34.0 DMLaw 51.5 21.4 22.4 25.2 18.2 39.0 50.7 19.4 52.6 29.8 33.0 RegMix 53.1 22.1 22.2 25.4 19.0 39.1 53.5 18.4 60.7 30.0 34.4 Influence functions MATES 52.6 21.8 22.6 26.7 20.4 40.9 53.7 19.7 57.6 31.8 34.8 Multi-Agent Collaboration (Ours) 65.8 31.5 23.0 26.6 24.6 39.9 54.1 20.1 60.4 30.5 37.7
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random sample Uniform-30B 54.5 21.9 22.4 25.6 19.2 39.7 54.2 19.7 53.2 30.8 34.1 Uniform-60B 59.1 26.0 22.4 26.9 21.6 42.1 54.3 21.0 55.7 32.4 36.2 Perplexity-based data selection PPL 49.2 21.2 22.5 24.9 14.6 36.7 49.8 20.6 60.6 23.3 32.4 Classifier-based data selection QuRating-Facts 60.5 25.4 23.4 26.4 20.2 40.3 51.9 19.0 58.0 23.3 34.8 QuRating-Req 59.9 26.4 22.8 25.6 21.8 40.1 53.7 19.6 56.9 22.3 34.9 QuRating-Writing 57.3 25.3 22.6 25.0 21.4 41.6 53.5 19.6 49.5 22.1 33.8 QuRating-Edu 60.8 26.5 22.5 26.5 20.2 41.4 54.6 21.1 55.5 22.7 35.2 FineWeb-Edu 56.6 24.9 22.6 25.8 19.8 39.4 51.2 19.7 55.9 30.9 34.7 DSIR-Book 49.7 21.1 22.1 25.6 19.8 41.7 54.1 18.3 55.6 22.9 33.1 DSIR-Wiki 53.6 22.3 23.0 25.3 17.6 36.7 52.2 20.4 60.2 22.6 33.4 Domain mixing methods DOGE 53.0 21.8 22.0 26.3 17.2 40.1 51.7 18.8 58.5 30.1 33.9 DoReMi 52.7 22.2 22.4 25.5 16.2 39.3 51.9 20.9 60.0 31.0 34.2 DMLaw 52.4 21.4 23.0 25.7 17.2 39.2 50.6 19.2 51.4 29.9 33.0 RegMix 53.5 24.0 21.2 25.0 19.6 41.0 53.2 19.0 61.3 30.2 34.8 Influence functions MATES 53.6 21.6 22.6 26.1 20.4 41.7 53.1 20.4 60.1 32.0 35.2 Multi-Agent Collaboration (Ours) 64.9 31.1 22.4 26.3 23.6 39.0 53.1 20.4 60.4 30.7 37.2
@@
A.4 Implementation Details of our Methods
@@
A.5 Details of Influence Changes During Different Pretraining Stages
@@
Appendix B Data Distribution Analysis of the SlimPajama Dataset

### v3

@@ A.7.1 Details of Pretraining Models Architecture
Hyperparameter 370M Model Value 1.3B Model Value 3.6B Model Value 8B Model
Vocabulary Size 32,000 32,000 128,256 128,256
MLP Ratio 8/3 8/3 8/3 3.5
Hidden Dimension Size 2048 1024 3072 4096
Number of Layers 24 24 28 32
Number of Attention Heads 16 8 24 32
Number of KV Attention Heads 16 8 8 8
RoPE Base 10,000 10,000 500,000 500,000
Maximum Context Window Length 1024 1024 1024 1024
Number of Parameters 373,867,520 (370M) 1,345,423,360 (1.3B) 3,606,752,256 (3.6B) 8,030,261,248 (8B)
A.7.2 Details of Baseline Method Implementation
@@
A.7.3 Details actor Weight Initialization
For the domain actor, we use the document’s meta-information, label the data with domain information and save this into the domain actor’s memory, where the domain $\text{Domain}(x_{i})$ belongs to one of ArXiv, Book, Wikipedia, CommonCrawl, GitHub, StackExchange, C4.
For quality actor, we adopt the FineWeb-Edu quality scorer (Lozhkov et al., 2024), which is fine-tuned as a BERT-like regression model Merrick et al. (2024) using Llama3-70B-Instruct annotated 500k examples. This will give out a successive quality score $\text{Quality}(x_{i})\in\mathcal{R}^{[0,5]}$ with higher score represent higher quality. We then map this score into five quality intervals $\{I_{j}\}_{j=1}^{5}$, as
$\begin{split}\text{Quality}(x_{i})\mapsto I_{j}=\begin{cases}[j-1,j),&\text{if}\text{Quality}(x_{i})\in[j-1,j),\\ &\quad j=1,2,3,4\\ [4,5],&\text{if }\text{Quality}(x_{i})\in[4,5],\\ &\quad j=5\end{cases}\end{split}$
We store the quality interval corresponding to each data point in the quality actor’s memory.
For the topic actor, due to the absence of a suitable pretrained model for topic classification and labeling, we designed a classification schema using 1.44 billion documents collected by the Common Crawl project (Project, 2007) and fine-tuned a BERT-like regression model on 500k GPT-4o annotated samples, the overall pipeline is depicted in Figure 4. Further details on the topic classification approach and BERT model training are provided in §A.2. Using this topic classifier, we categorize each document into one of 13 topics: Activity, Education, Entertainment, Finance, Health, Business and Industrial, Infrastructure, Literature and Art, Nature, Others, Law and Government, Networking, Technology, and store the topic information in the topic actor’s memory.
We employ an actor weight initialization technique within the RegMix (Liu et al., 2024) framework, which is crucial for the effective training of proxy models. Our dataset is organized into three distinct categories: domain, quality, and topic. For each category, we initialize the data weights based on the original proportions across 512 configurations and subsequently train a TinyLlama-1M with 1 billion tokens as a proxy model for each configuration. We evaluate this model on previously unseen data mixtures, specifically using validation set loss, following RegMix, for assessment. We then fit a regression model based on the performance results of the 512 proxy models to predict the optimal data mixture for training large-scale LMs. The results of the LightGBM regression analysis and Spearman correlation of the loss prediction performance are presented in 7.
@@ A.7.3 Details actor Weight Initialization
Figure 7: We present the results of the LightGBM regression analysis and Spearman correlation regarding the loss prediction performance and the weights of each candidate data-(a) Topic, (b) Quality Interval and (c) Domain after mixture across all categories.
@@
A.8 Full Experimental Results
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random sample Uniform-30B 54.3 23.4 22.3 23.9 18.6 39.8 52.8 19.2 55.4 30.0 34.0 Uniform-60B 55.2 24.6 22.5 23.4 21.0 39.7 51.9 19.5 59.8 33.1 35.1 Perplexity-based data selection PPL 49.3 20.1 22.4 23.6 16.2 36.0 48.1 18.8 61.4 29.3 32.5 Classifier-based data selection QuRating-Facts 56.1 23.3 22.4 24.8 21.6 39.2 54.1 19.9 61.5 31.6 35.5 QuRating-Req 54.9 24.4 23.2 25.2 21.4 38.1 54.5 20.6 61.6 31.3 35.5 QuRating-Writing 53.6 23.2 23.4 23.2 21.0 38.1 52.8 19.7 59.4 31.6 34.6 QuRating-Edu 57.0 24.4 22.0 25.0 20.4 40.3 53.7 20.2 60.1 32.2 35.5 FineWeb-Edu 53.8 23.4 21.8 23.9 19.8 39.2 51.7 20.8 59.7 32.0 34.6 DSIR-Book 45.4 20.8 22.0 23.0 18.8 39.9 54.6 19.7 58.3 30.8 33.3 DSIR-Wiki 50.6 21.1 21.6 23.0 19.2 36.6 53.0 19.8 60.5 29.2 33.5 Domain mixing methods DOGE 49.4 21.8 22.5 23.0 18.0 38.0 52.7 19.9 60.0 30.0 33.5 DoReMi 50.1 20.2 22.5 23.7 17.8 38.7 52.8 19.7 58.6 30.8 33.5 DMLaw 49.6 21.9 23.2 23.6 17.8 38.6 51.8 20.1 60.4 29.0 33.6 RegMix 50.0 22.3 22.1 22.9 18.8 38.0 52.8 19.9 58.9 31.2 33.7 Influence functions MATES 50.0 21.4 22.7 25.3 19.0 39.8 53.6 21.3 59.9 32.1 34.5 Multi-actor Collaboration (Ours) 61.1 28.2 22.6 26.0 24.4 38.2 54.2 19.5 61.0 29.8 36.5
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random sample Uniform-30B 54.6 23.0 22.1 24.9 18.8 40.3 52.9 21.5 53.0 29.8 34.1 Uniform-60B 58.8 25.5 23.0 27.2 20.0 41.8 53.6 19.6 56.9 32.7 35.9 Perplexity-based data selection PPL 50.6 21.3 22.7 25.2 15.6 37.7 48.9 20.1 61.5 22.3 32.6 Classifier-based data selection QuRating-Facts 59.5 25.7 22.6 25.9 19.8 40.2 54.6 19.2 60.8 24.8 35.3 QuRating-Req 59.3 25.9 22.7 26.1 19.6 39.7 53.7 20.5 58.5 22.7 34.9 QuRating-Writing 56.9 25.7 23.1 26.0 20.4 41.1 53.6 20.2 51.4 22.6 34.1 QuRating-Edu 60.8 26.5 22.5 26.7 20.2 41.4 54.6 20.6 55.5 22.7 35.1 FineWeb-Edu 56.2 25.7 22.3 26.2 20.6 40.1 50.5 19.7 56.6 31.4 34.9 DSIR-Book 48.7 21.0 22.6 25.6 18.6 42.5 53.7 19.5 57.9 22.9 33.3 DSIR-Wiki 53.2 22.4 22.6 25.3 17.6 37.1 52.7 21.4 61.6 24.2 33.8 Domain mixing methods DOGE 52.4 21.9 22.4 27.0 17.4 39.9 52.0 18.2 57.8 29.8 33.9 DoReMi 53.2 21.4 22.2 24.7 18.2 38.4 50.9 20.6 59.7 31.1 34.0 DMLaw 51.5 21.4 22.4 25.2 18.2 39.0 50.7 19.4 52.6 29.8 33.0 RegMix 53.1 22.1 22.2 25.4 19.0 39.1 53.5 18.4 60.7 30.0 34.4 Influence functions MATES 52.6 21.8 22.6 26.7 20.4 40.9 53.7 19.7 57.6 31.8 34.8 Multi-actor Collaboration (Ours) 65.8 31.5 23.0 26.6 24.6 39.9 54.1 20.1 60.4 30.5 37.7
@@
Problem Solving Commonsense Reasoning Reading Compreh. Selection Method ARC-E ARC-C MathQA MMLU O.B.QA SIQA W.G. C.S.QA BoolQ RACE Average Random sample Uniform-30B 54.5 21.9 22.4 25.6 19.2 39.7 54.2 19.7 53.2 30.8 34.1 Uniform-60B 59.1 26.0 22.4 26.9 21.6 42.1 54.3 21.0 55.7 32.4 36.2 Perplexity-based data selection PPL 49.2 21.2 22.5 24.9 14.6 36.7 49.8 20.6 60.6 23.3 32.4 Classifier-based data selection QuRating-Facts 60.5 25.4 23.4 26.4 20.2 40.3 51.9 19.0 58.0 23.3 34.8 QuRating-Req 59.9 26.4 22.8 25.6 21.8 40.1 53.7 19.6 56.9 22.3 34.9 QuRating-Writing 57.3 25.3 22.6 25.0 21.4 41.6 53.5 19.6 49.5 22.1 33.8 QuRating-Edu 60.8 26.5 22.5 26.5 20.2 41.4 54.6 21.1 55.5 22.7 35.2 FineWeb-Edu 56.6 24.9 22.6 25.8 19.8 39.4 51.2 19.7 55.9 30.9 34.7 DSIR-Book 49.7 21.1 22.1 25.6 19.8 41.7 54.1 18.3 55.6 22.9 33.1 DSIR-Wiki 53.6 22.3 23.0 25.3 17.6 36.7 52.2 20.4 60.2 22.6 33.4 Domain mixing methods DOGE 53.0 21.8 22.0 26.3 17.2 40.1 51.7 18.8 58.5 30.1 33.9 DoReMi 52.7 22.2 22.4 25.5 16.2 39.3 51.9 20.9 60.0 31.0 34.2 DMLaw 52.4 21.4 23.0 25.7 17.2 39.2 50.6 19.2 51.4 29.9 33.0 RegMix 53.5 24.0 21.2 25.0 19.6 41.0 53.2 19.0 61.3 30.2 34.8 Influence functions MATES 53.6 21.6 22.6 26.1 20.4 41.7 53.1 20.4 60.1 32.0 35.2 Multi-actor Collaboration (Ours) 64.9 31.1 22.4 26.3 23.6 39.0 53.1 20.4 60.4 30.7 37.2
@@
A.9 Implementation Details of our Methods
@@
A.10 Details of Influence Changes During Different Pretraining Stages
@@
A.11 Data Distribution Analysis of the SlimPajama Dataset